9 16/64 Reduce to simplest form
Answer by sdmmadam@yahoo.com(530)

9 16/64
Is it problem 9) and the fraction is 16/64
Then 16/64 = 16/(16X4) = 1/4 (Cancelling 16)
OR is it 916/64
916/64
= 2X458/2X32 (Cancelling 2)
= 458/32
=2X229/2X16 (Cancelling 2)
=229/16
Now the nr and the dr have only the factor 1 as common. This is the sign to see that the fraction is in the simplest form.